Two hockey pucks, labeled A and B, are initially at rest on a smooth ice surface and are separated by a distance of 18.0 m . Simultaneously, each puck is given a quick push, and they begin to slide directly toward each other. Puck A moves with a speed of 1.50 m/s , and puck B moves with a speed of 2.50 m/s .

(Since (vor n has a negative value and (o)A has a positive value, the pucks are moving in opposite directions) Review Learning Goal: To practice Problem-Solving Strategy for general problems Part D Two hockey pucks, labeled A and B, are initially at rest on a smooth ice surface and are separated by a distance of 18.0 m. Simultaneously, each puck is given a quick push, and they begin to slide directly toward each other. Puck A moves with a speed of 1.50 m/s, and puck B moves with a speed of 2.50 m/s What is 1, the distance that puck A covers prior to the collision? Express your answer in meters to three significant figures. Do not include units in your answer. View Available Hint(s) What is the distance covered by puck A by the time the two pucks collide? Im Check that you are using commas appropriately in your answer No credit lost. Try again. Submit us Ans uest Answer Part E Complete previous part(s) Provide Feedback
Answer #1

speed of A = 1.5 m/s

B = 2.5 m/s

A and B approach other

velocity of approach = 1.5 + 2.5 = 4.0 m/s ( velocity of each puck relative to the other.

time of collision = 18/4 = 4.5 s ( they are separated by 18 m)

each puck travels for 4.5 s before collision

distanced moved by A = 1.5 *4.5 = 6.75 m